# REA-RL: Reflection-Aware Online Reinforcement Learning for Efficient Large Reasoning Models

### v1

@@ 5.2 Main Result
Budget: 8k Original 91.58 1001312 88.40 1003389 77.40 1003270 77.50 1004746 51.70 1005467 40.83 1006815 71.24 1004166
SFT 89.92 64.63 89.20 69.25 77.92 72.91 82.81 72.67 52.44 82.06 41.25 89.98 72.26 75.25
RPO 90.14 51.22 89.80 53.35 83.12 55.20 80.62 59.54 58.52 61.24 41.67 74.78 73.98 59.22
GRPO 92.80 117.99 89.00 100.74 79.22 101.31 79.69 96.23 51.70 100.00 39.17 98.88 71.93 102.52
GRPO $\text{R}_{\text{Len}}$ 85.97 32.01 87.60 59.04 70.91 58.84 85.62 71.43 55.56 76.59 45.42 87.42 71.85 64.22
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 37.80 91.20 48.66 78.44 52.23 86.56 59.65 57.19 70.37 41.67 84.15 74.05 58.81
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 67.07 91.40 69.52 80.52 72.54 85.62 75.64 53.93 83.04 47.92 91.37 75.35 76.53
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 53.12 89.60 61.79 81.30 61.19 85.62 71.39 55.11 77.81 42.92 88.45 74.09 73.86
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 60.15 1007549 48.33 10010460 77.01 1005478
SFT 89.99 68.75 90.60 69.25 79.48 74.37 88.44 72.11 57.63 84.70 48.75 89.68 75.82 76.48
RPO 90.14 50.00 89.80 46.90 83.12 47.77 82.19 50.75 58.37 45.21 42.92 51.41 74.42 48.67
GRPO 92.87 116.52 93.40 99.20 82.34 100.26 87.19 97.31 60.44 96.78 50.42 97.18 77.78 101.21
GRPO $\text{R}_{\text{Len}}$ 85.97 31.25 88.40 56.43 72.21 55.88 87.81 65.65 59.56 69.20 50.00 76.96 73.99 59.23
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 36.90 92.40 45.11 79.74 47.50 88.12 56.04 60.74 63.52 47.92 75.82 76.36 54.15
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 65.48 92.80 66.71 81.82 68.27 88.75 72.17 58.81 77.37 54.58 86.21 78.25 72.70
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 52.08 91.00 60.78 82.08 55.85 89.38 67.76 59.11 73.11 51.25 81.09 77.13 65.11
Table 2: Main results of our proposed methods. Most abbreviations align with Table 1. Baseline definitions are in §5.1. “GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$” represents the addition of the reflection model, “GRPO $\text{R}_{\text{RLen+Reflect}}$” represents the addition of the reward optimization, and “GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$” represents the combination of both optimizations. “Budget” on the left is the max tokens allowed per question. Generation is forcibly stopped if this limit is exceeded.
Improved performance compared to offline method.
The results are in Table 2. The final three rows showcase the performance of the reflection model, reward refinement, and their combined approach. While RPO shares similar ideas with the sequential revision, the primary distinction lies in the offline dataset. RPO achieves comparable performance at an 8k budget but significantly degrades on the three most challenging datasets at a 16k budget. In contrast, our method demonstrates better performance, highlighting the advantage of online training over offline training.
GRPO, using only an accuracy reward, shows no significant performance improvement, suggesting that gains are not solely due to more training. Adding a length reward greatly reduces inference costs but severely hurts performance. Our method significantly improves the probability of solving problems within an 8k token budget (3.9% to 5.8% on average) and effectively shortens responses. With a sufficient 16k token budget, our approach reduces response length by 27% to 46% while maintaining accuracy.
Reflection model and reflection reward target distinct dimensions.
The reflection model is more effective in reducing response length, whereas reflection reward contributes more to accuracy enhancement. By integrating both strategies, a balanced outcome could be achieved, yielding a 35% efficiency improvement without compromising performance. However, the improvement under an 8k budget remains limited. This can be attributed to the advantage of shorter generations under constrained budget, and the reflection model’s inherent ability to foster reflection, as analyzed in §5.4. The effectiveness of $\text{R}_{\text{RLen}}$ and $\text{R}_{\text{Reflect}}$ is further separately verified in Appendix B.2.
5.3 Reflection Model Analysis
Method GSM8K Math500 Gaokao23 Amc23 Olympiad Aime24 Average
@@ 5.3 Reflection Model Analysis
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 60.15 1007549 48.33 10010460 77.01 1005478
32B Revise 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 60.74 83.06 50.83 93.59 77.30 76.88
 + Gold 88.78 54.09 92.00 57.80 79.48 59.71 87.19 67.71 60.15 73.48 51.67 92.31 76.54 67.52
7B Revise Weak 90.75 84.82 92.40 86.64 82.86 86.71 88.75 88.90 62.22 90.63 50.00 93.10 77.83 88.47
Fixed Trunc (88.47) 89.84 84.23 92.00 86.75 81.04 86.92 88.12 89.30 61.19 91.27 47.50 93.55 76.61 88.67
7B Revise Normal 90.22 79.02 91.80 81.09 82.86 81.82 89.06 83.80 61.78 87.55 49.17 91.75 77.48 84.17
Fixed Trunc (84.17) 89.23 78.50 91.40 82.02 80.52 82.51 86.25 85.14 59.85 88.04 46.67 92.48 75.65 84.78
7B Revise Strong 90.07 72.99 92.00 75.65 82.08 77.75 87.81 77.16 60.59 80.99 48.75 89.04 76.88 78.93
Fixed Trunc (78.93) 88.32 72.77 90.00 76.73 79.22 78.76 80.94 80.02 58.96 82.08 45.83 90.12 73.88 80.08
GRPO 92.87 116.52 93.40 99.20 82.34 100.26 87.19 97.31 60.44 96.78 50.42 97.18 77.78 101.21
GRPO Gen8 92.42 103.35 92.00 97.82 82.34 100.24 90.00 92.74 58.07 98.68 50.00 93.60 77.47 97.74
GRPO $\text{R}_{\text{Len}}$ 85.97 31.25 88.40 56.43 72.21 55.88 87.81 65.65 59.56 69.20 50.00 76.96 73.99 59.23
GRPO $\text{R}_{\text{Len}}$ Gen8 85.52 25.74 88.20 51.68 72.21 56.78 83.12 63.78 57.78 71.53 53.33 82.51 73.36 58.67
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 36.90 92.40 45.11 79.74 47.50 88.12 56.04 60.74 63.52 47.92 75.82 76.36 54.15
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 52.08 91.00 60.78 82.08 55.85 89.38 67.76 59.11 73.11 51.25 81.09 77.13 65.11
Table 3: Results of our proposed reflection model. “32B Revise” refers to the “Model Revise” results using a 32B LLM from Table 1. “7B Revise” uses our trained 7B revision model. “Fixed Trunc” indicates the use of a fixed truncation strategy with the same ratio as our corresponding method. Truncation strengths are defined in §5.3. The last three rows are copied from Table 2.
Experimental setup.
We follow the setup in §3.2 to verify our reflection model. We define three revision strengths: Normal truncates at the first identified correct answer; Weak truncates at the second such position; and Strong truncates before the first position where the truncation probability exceeds 0.25. If no such position exists, the Normal position is used. Fixed Trunc represents truncation at the closest sentence-ending position corresponding to the truncation ratio achieved by our method, and then forces the LRM to generate the final answer, as defined in §3.2.
Comparable performance to 32B at a significantly lower cost.
Results are shown in Table 3. Our reflection model does not incorporate the gold answer as input. In comparison to the 32B model revised without gold input, our Strong strategy achieves a comparable compression ratio while exhibiting minimal performance degradation, demonstrating that our model achieves comparable results at a substantially reduced cost. Furthermore, our method also outperforms fixed truncation, especially when the truncation ratio is large. Finally, we observe that under certain evaluations, the accuracy actually increases after revision. This indicates that the model may have already arrived at the correct answer but fails to explicitly state this conclusion under the 16k token budget.
REA-RL outperforms inference with reflection model.
To ensure revision accuracy, we employ the Normal strategy during training. However, REA-RL with the reflection model achieves a higher compression ratio than the Strong strategy while maintaining comparable performance, despite the increased inference cost associated with the generate-then-revise approach of Strong strategy. This improvement can be attributed to the length reward and the online training, which iteratively train the model and provide further guidance for shortening already concise responses.
Online revision as an efficient scaling strategy.
The reflection model provides revised responses, which doubles the dataset size. To evaluate whether other methods could yield similar benefits, we extend the GRPO baseline generation to eight paths, i.e., “Gen8”, to verify whether parallel sampling scaling is more effective. However, Gen8 indicates no consistent improvement over parallel sampling of four paths, while our combination with sequential revision, under the same reward $\text{R}_{\text{Len}}$, demonstrates superior performance and truncation ratios, thereby establishing the enhanced efficacy of our scaling method compared to mere parallel sampling.
Regarding training time, we optimize the implementation by deferring the update of the vllm inference model, which enables parallel execution of training and data generation, resulting in approximately a twofold speedup. On 3 NVIDIA A800 80G GPUs, GRPO with sampling 4 paths requires 80 hours, while sampling 8 paths requires 110 hours. Our method, incorporating the reflection model, required 120 hours. Given the significant performance improvement, this additional computational cost is deemed acceptable. If training efficiency is a primary concern, utilizing the reflection reward alone also requires only 80 hours and achieves improved performance.
@@ 5.4 Reflection Ability Analysis for the Trained Model
Method GSM8K Math500 Gaokao23 Amc23 Olympiad Aime24 Average
[…]
Budget: 16k Original 91.66 105.48 92.00 107.75 81.82 90.94 88.12 102.72 60.15 80.09 48.33 84.82 77.01 95.30
SFT 89.99 87.73 90.60 104.43 79.48 79.97 88.44 90.00 57.63 77.31 48.75 82.30 75.82 86.96
RPO 90.14 156.68 89.80 146.65 83.12 130.13 82.19 127.03 58.37 101.43 42.92 100.61 74.42 127.09
GRPO 92.87 97.50 93.40 107.52 82.34 88.69 87.19 98.06 60.44 79.10 50.42 82.28 77.78 92.19
GRPO $\text{R}_{\text{Len}}$ 85.97 813.96 88.40 128.32 72.21 114.74 87.81 120.09 59.56 93.71 50.00 99.07 73.99 228.32
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 194.97 92.40 135.54 79.74 120.05 88.12 107.91 60.74 87.01 47.92 89.34 76.36 122.47
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 141.63 92.80 118.61 81.82 99.24 88.75 106.08 58.81 78.08 54.58 86.13 78.25 104.96
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 150.87 91.00 126.51 82.08 109.91 89.38 105.89 59.11 89.82 51.25 90.25 77.13 112.21
Table 4: Reflection density of our proposed methods and baselines. Most abbreviations align with Table 2. “Reflect” represents the average number of tokens between each reflective token.
Figure 3: Accuracy and generation length changes during training on GSM8k. The x-axis represents the training steps, with a maximum of 1800 steps. The left plot shows the accuracy of question answering, and the right plot shows the average token consumption per answer. We present results for the GRPO baseline using only a length reward and our three proposed methods.
@@
We calculate the average number of tokens between reflective tokens in the responses, and the results are in Table 4. Introducing only a length reward in online RL leads to a significant decrease in the frequency of reflective tokens in simple problems. However, for challenging problems, the model retains its reflection capability, as reflection is crucial for obtaining the accuracy reward. A case is in the middle of Figure 1, where the solution only performs planning in the “think” part and solves the problem without any reflection. Due to the minimal token consumption, an error occurs during planning. In contrast, our method preserves the style of R1, reasoning for sufficient tokens, and performing reflection after reasoning. To provide further analysis, we further illustrate the training dynamics of GSM8k in Figure 3, demonstrating that it sacrifices accuracy to achieve token length reduction, showing similar trends between accuracy and response length.
@@
Across the three easier datasets, our three approaches demonstrate an average increase of 49% in reflection probability and 7.5% performance improvement compared to GRPO $\text{R}_{\text{Len}}$, demonstrating that our methods effectively mitigate the issue of insufficient reflection. As illustrated in the training dynamics, the reflection reward facilitates a gradual improvement in efficiency with minimal change in accuracy. Besides, although using the reflection model itself initially leads to performance degradation similar to the baseline, the accuracy reward penalizes overly short responses that result in low accuracy, allowing the accuracy to recover. During the recovery, the reflection model can provide concise and more accurate training data, aiding the model in recovering performance without compromising efficiency. Finally, given that the reflection model inherently promotes reflection to some extent, its combination with the reflection reward can only achieve a trade-off rather than a substantial further improvement.
Mitigating overthinking on simple problems.
LRM often exhibits overthinking on easier problems. Our method appropriately reduces reflection on easier problems while maintaining it on more difficult ones. While our approach increases the reflection frequency on simpler problems compared to $\text{R}_{\text{Len}}$, it remains lower than that of the original model. Specifically, on the three simplest datasets, we reduce the reflection frequency by 31%, and 45% efficiency enhancement. Whereas on challenging ones, the reduction of reflection is only 5%, and 27% efficiency enhancement. This demonstrates that our approach achieves a favorable balance, mitigating overthinking on simple problems while preserving reflection capabilities. Furthermore, we provide additional evidence demonstrating that our method reduces overthinking while preserving reflection capabilities in Appendix B.3.
@@ 6 Conclusion and Limitations
To enhance the inference efficiency of LRM without compromising model performance, we propose REflection-Aware online Reinforcement Learning, REA-RL, to achieve improved online scaling and better performance retention. Specifically, we introduce a reflection model for efficient scaling, offering sequential revision to augment parallel sampling data generation for online RL. Furthermore, we introduce a reflection reward to better maintain model performance. Experiments show that REA-RL reduces inference cost by 35% with no performance loss. Analysis proves that our method is effective by maintaining reflection frequency for challenging problems while appropriately reducing it for simpler ones, thus balancing performance and efficiency.
Our paper has the following limitations. First, our approach is only validated on distilled 7B LRMs. Due to the large model size and long training time, we do not perform validation on LRMs pre-trained from scratch, which aligns with prior work [46, 15, 32, 2]. Second, the detection method proposed in §3 is based on LLMs. While it outperforms fixed truncation, it cannot guarantee the complete elimination of overthinking. Nevertheless, our primary goal is to use it to train the reflection model, and as long as it can reduce overthinking to a certain extent, it can effectively shorten the reasoning process during iterative training. Finally, our scaling method introduces approximately 10% additional computational cost compared to using parallel scaling only. This is due to the use of sequential scaling, which results in poorer parallelism and longer runtime. However, considering our significant improvement and the cost-free improvement of reflection reward, this is acceptable.

### v2

@@ 5.3 Reflection Model Analysis
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 48.33 10010460 80.39 1005064
7B Revise 86.50 63.32 89.20 64.60 77.40 66.16 82.81 73.17 42.92 80.98 75.77 69.65
  + Gold 85.90 55.36 85.60 56.59 76.62 55.96 79.06 61.25 42.92 76.58 74.02 61.15
32B Revise 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 50.83 93.59 80.61 75.65
  + Gold 88.78 54.09 92.00 57.80 79.48 59.71 87.19 67.71 51.67 92.31 79.82 66.32
$\text{M}_{\text{Reflect}}$ Weak 90.75 84.82 92.40 86.64 82.86 86.71 88.75 88.90 50.00 93.10 80.95 88.03
Fixed Trunc (88.47) 89.84 84.23 92.00 86.75 81.04 86.92 88.12 89.30 47.50 93.55 79.70 88.15
$\text{M}_{\text{Reflect}}$ Normal 90.22 79.02 91.80 81.09 82.86 81.82 89.06 83.80 49.17 91.75 80.62 83.50
Fixed Trunc (84.17) 89.23 78.50 91.40 82.02 80.52 82.51 86.25 85.14 46.67 92.48 78.81 84.13
$\text{M}_{\text{Reflect}}$ Strong 90.07 72.99 92.00 75.65 82.08 77.75 87.81 77.16 48.75 89.04 80.14 78.52
Fixed Trunc (78.93) 88.32 72.77 90.00 76.73 79.22 78.76 80.94 80.02 45.83 90.12 76.86 79.68
GRPO 92.87 116.52 93.40 99.20 82.34 100.26 87.19 97.31 50.42 97.18 81.24 102.09
GRPO Gen8 92.42 103.35 92.00 97.82 82.34 100.24 90.00 92.74 50.00 93.60 81.35 97.55
GRPO $\text{R}_{\text{Len}}$ 85.97 31.25 88.40 56.43 72.21 55.88 87.81 65.65 50.00 76.96 76.88 57.23
GRPO $\text{R}_{\text{Len}}$ Gen8 85.52 25.74 88.20 51.68 72.21 56.78 83.12 63.78 53.33 82.51 76.48 56.10
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 36.90 92.40 45.11 79.74 47.50 88.12 56.04 47.92 75.82 79.48 52.27
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 65.48 92.80 66.71 81.82 68.27 88.75 72.17 54.58 86.21 82.13 71.77
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 52.08 91.00 60.78 82.08 55.85 89.38 67.76 51.25 81.09 80.74 63.51
Table 3: Results of our proposed reflection model. “7B Revise” and “32B Revise” refer to the two-step revision method introduced in §3, using Qwen-7B and Qwen-32B without training. $\text{M}_{\text{Reflect}}$” uses our 7B reflection model with one-step revision. “Fixed Trunc” denotes a fixed truncation strategy with the same ratio as our method. Truncation strengths are defined in §5.3.
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average
@@
Budget: 16k Original 91.66 105.48 92.00 107.75 81.82 90.94 88.12 102.72 48.33 84.82 80.39 98.34
SFT 89.99 87.73 90.60 104.43 79.48 79.97 88.44 90.00 48.75 82.30 79.45 88.89
RPO 90.14 156.68 89.80 146.65 83.12 130.13 82.19 127.03 42.92 100.61 77.63 132.22
GRPO 92.87 97.50 93.40 107.52 82.34 88.69 87.19 98.06 50.42 82.28 81.24 94.81
GRPO $\text{R}_{\text{Len}}$ 85.97 813.96 88.40 128.32 72.21 114.74 87.81 120.09 50.00 99.07 76.88 255.24
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 194.97 92.40 135.54 79.74 120.05 88.12 107.91 47.92 89.34 79.48 129.56
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 141.63 92.80 118.61 81.82 99.24 88.75 106.08 54.58 86.13 82.13 110.34
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 150.87 91.00 126.51 82.08 109.91 89.38 105.89 51.25 90.25 80.74 116.69
Table 4: Reflection density of REA-RL and baselines. “Reflect” represents the average number of tokens between each reflective token, i.e., a smaller value indicates more frequent reflection.
@@ 5.4 Reflection Ability Analysis for the Trained Model
We calculate the average number of tokens between reflective tokens in the responses, and the results are in Table 4. Introducing only a length reward in online RL leads to a severe decrease in the frequency of reflective tokens on easy problems. However, for challenging problems, the model retains its reflection capability, as reflection is crucial for obtaining the accuracy reward. An example is shown in the middle of Figure 1, where the solution only performs planning in the “think” part and solves the problem without any reflection. Due to the minimal token consumption, an error occurs during planning. In contrast, our method preserves the style of R1 by reasoning with sufficient tokens, and performing reflection after reasoning.
@@
Across the three easier datasets, our three approaches yield an average 47% increase in reflection probability and a 5% performance improvement over GRPO $\text{R}_{\text{Len}}$, demonstrating that REA-RL mitigates non-reflective behavior.
Mitigating overthinking on easy problems.
LRMs often overthink on easy problems. Our method appropriately reduces reflection on easy problems while maintaining it on difficult ones. While our approach increases the reflection frequency on easier problems compared to $\text{R}_{\text{Len}}$, it remains lower than that of the original model. Specifically, on the three easiest datasets, we reduce the reflection frequency by 22%, with a 45% efficiency enhancement. Whereas on challenging ones, the reduction in reflection is only 4%, with a 27% efficiency enhancement. This demonstrates that our approach achieves a favorable balance, mitigating overthinking on easy problems while preserving reflection capabilities. Furthermore, we provide additional evidence demonstrating that our method reduces overthinking while preserving reflection capabilities in Appendix B.6.
Training dynamics analysis.
To provide further analysis, we illustrate the training dynamics in Figure 3. For GRPO $\text{R}_{\text{Len}}$, after 1,000 steps, when the length falls below 70% of its original value, its performance begins to drop, indicating that the model begins to remove meaningful steps. However, the reflection reward enables a gradual improvement in efficiency with minimal change in accuracy. Besides, the reflection model greatly accelerates length reduction. Although it initially causes severe performance degradation, the targeted penalties that it imposes on overthinking tokens avoid deleting valid reflection tokens, halt further length shortening, and gradually recover performance. After 1,000 steps, it consistently outperforms the baseline in both accuracy and efficiency. Finally, the reflection model’s goal remains to penalize overthinking, which is inherently at odds with the reflection reward that indiscriminately encourages reflection. Therefore, their combination achieves a trade-off between efficiency and accuracy, yielding a balance rather than a synergistic gain.
Figure 3: Changes in accuracy and generation length during training on five test sets on average. The x-axis represents the training steps. The left plot shows the average accuracy, and the right plot shows the average token consumption per answer. Abbreviations are aligned with Table 2.
@@ 6 Conclusion and Limitations
To enhance the inference efficiency of LRMs without compromising model performance, we propose REflection-Aware online Reinforcement Learning, REA-RL, to achieve improved online scaling and better performance retention. Specifically, we introduce a reflection model for efficient scaling, offering sequential revision to augment parallel sampling data generation for online RL. Furthermore, we introduce a reflection reward to better maintain model performance. Experiments show that REA-RL reduces token cost by 36% with no performance loss. Analysis shows that our method is effective by maintaining reflection frequency for challenging problems while appropriately reducing it for easier ones, thus balancing performance and efficiency.
Our paper has the following limitations. First, our approach is only validated on distilled 7B LRMs. Due to the large model size and long training time, we do not perform validation on LRMs pre-trained from scratch, which aligns with prior work (Zhang and Zuo, 2025; Huang et al., 2025; Qu et al., 2025b). Second, the detection method proposed in §3 is based on LLMs. While it outperforms fixed truncation, it cannot guarantee the complete elimination of overthinking. Nevertheless, our goal is to train the reflection model, and as long as it can reduce overthinking to a certain extent, it can effectively shorten the reasoning process during iterative training. Finally, our scaling method introduces approximately 10% additional cost compared to using parallel scaling only. This is due to the use of sequential scaling, which results in poorer parallelism for vllm. However, considering our significant improvement and the cost-free improvement of the reflection reward, this is acceptable.
Ethics Statement
Our work adheres to the ICLR Code of Ethics and uses publicly available datasets for reproducibility. LLMs may exhibit racial and gender biases, so we strongly recommend users assess potential biases before applying the models in specific contexts. Additionally, due to the difficulty of controlling LLM outputs, users should be cautious of issues arising from hallucinations.
Reproducibility Statement
We make our code, configuration files, and evaluation scripts available at the anonymous repository linked in the abstract (https://github.com/hexuandeng/REA-RL ). All hyperparameters required to reproduce our method are provided in Appendix B.2. The required hardware and runtime are reported in §5.3. The use of large language models is discussed in Appendix B.7.
Acknowledgments
This work was supported in part by the Guangdong S&T Program (Grant No. 2024B0101050003), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011491), and Shenzhen Science and Technology Program (Grant Nos. ZDSYS20230626091203008,KJZD202310230947 00001,KQTD2024072910215406).